Question 6 of 17 (1 point) Evalu Evaluate the function for the indicated values. f(x) = 7x2 – 2x + 6 (2) is $(-3) is
Answer #1

Ans:

Given that

f(x)=7x^2-2x+6

when x=2

f(2)=7*2^2-2*2+6

=28-4+6

=30

when x=-3

f(-3)=7*(-3)^2-2*(-3)+6

=7*9+2*3+6

=63+6+6

=75